Accumulate PhP 180,000 for 5 years at 13% compounded semi-annually. (Note: Accumulating an amount means finding the maturity value.)
Answer by mananth(16946)


Find Amount
Principal P = 180000
Amount= A
years=t 5.00
compounded 2 times a year n
Rate = 13.00 0.13 0.065
Amount = P*((n+r)/n)^n*t

Amount = = 180000 *( 1 + 0.065 )^ 2 * 5.00
Amount = 180000 *( 1 + 0.065 )^ 10
180000 *( 1.065 )^ 10.00
Amount = pHp337884.74